OpenStudy (anonymous):

3x² + 21x = 12, can someone explain to me how to find the value of x?

pooja195 (pooja195):

Well is this meant to be one of those quadratic equations?

OpenStudy (anonymous):

yes, its something ive been having troubles with.

OpenStudy (anonymous):

whoops, the question was mistyped, here is the equasion 3x² + 21x + 2 = 14

pooja195 (pooja195):

ok \[\huge~\rm~ax^2+bx+c=0 \] is this familiar?

OpenStudy (anonymous):

yeah.

pooja195 (pooja195):

Alright so make your given equation = 0 then list your abc values

OpenStudy (anonymous):

3x² + 21x + 2 = (0?) a = 3 b = 21 c = 2

pooja195 (pooja195):

We need to subtract 2-14=?

OpenStudy (anonymous):

3x² + 21x = 12?

pooja195 (pooja195):

No... it would -12 then you would do 3x^2+21x-12=0 Now list the abc values

OpenStudy (anonymous):

i think i figured it out now, thanks

pooja195 (pooja195):

what did you get?